# YOLO Evolution: A Comprehensive Benchmark and Architectural Review of YOLOv12, YOLO11, and Their Previous Versions

### v3

@@ IV-A1 Traffic Signs Dataset
Table VI presents a comparative analysis of the YOLO algorithms’ performance on the Traffic Signs dataset, evaluated based on accuracy, computational efficiency, and model size. The Traffic Signs dataset is a medium-sized dataset with varied object sizes, making it favorable for benchmarking. The results highlight the effectiveness of YOLO models in detecting traffic signs, demonstrating a range of precision. The highest mAP50-95 was 0.799, while the lowest recorded precision was 0.64. On the other hand, the highest mAP50 is 0.893 while the lowest is 0.722. The substantial gap between the mAP50 and mAP50-95 results suggests that the models have difficulties in uniformly handling traffic signs with different sizes at higher thresholds, reflecting areas for potential improvement in their detection algorithms.
@@
As illustrated in Figure 11, YOLOv5ul demonstrates the highest accuracy, achieving a mAP50 of 0.866 and a mAP50-95 of 0.799. This is followed by YOLO11m with a mAP50-95 of 0.795 and YOLO11l with a mAP50-95 of 0.794. In contrast, YOLOv10n exhibits the lowest precision, with a mAP50 of 0.722 and a mAP50-95 of 0.64, closely followed by YOLOv5un with a mAP50-95 of 0.665, as evidenced by the data points in Figure 11.
Figure 11: mAP50 and mAP50-95 YOLO results on traffic signs dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score.
Figure 12: Precision vs. Recall based on size on traffic signs dataset, with larger circles indicating larger model sizes.
Precision and Recall:
Figure 12 elucidates the trade-off between precision and recall taking the size of the models into consideration. Models such as YOLO11m, YOLO10l, YOLOv12m, YOLOv9m, YOLOv5ux, and YOLO11l exhibit high precision and recall, specifically with YOLO11m achieving a precision of 0.898 and a recall of 0.826 while having a size of 67.9Mb, and YOLOv12m achieving a precision of 0.868 and a recall of 0.811 with a significantly smaller size (39 Mb). In contrast, smaller models such as YOLOv10n (precision 0.722, recall 0.602), YOLOv8n (precision 0.749, recall 0.688), and YOLO11n (precision 0.768, recall 0.695) underperform in both metrics. This underscores the superior performance of larger models on the Traffic Signs dataset. Moreover, the high precision (0.883) and low recall (0.684) of YOLOv12n indicate a propensity for false negatives, while YOLOv3u’s high recall (0.849) and low precision (0.75) suggest a tendency for false positives.
Figure 13: Total processing time and GFLOPs count results on traffic signs dataset.
@@
In terms of computational efficiency, YOLOv10n is the most efficient, with a processing time of 2ms per image and a GFLOPs count of 8.3, as shown in Figure 13. YOLO11n closely trails this at 2.2ms with a 6.4 GFLOPs count, and YOLOv3u-tiny with a processing time of 2.4ms and a GFLOPs count of 19, making it relatively computationally inefficient compared to the other fast models. However, the data indicates that YOLOv9e, YOLOv9m, YOLOv9c, and YOLOv9s are the least efficient, with inference times of 16.1ms, 12.1ms, 11.6ms, and 11.1ms, and GFLOPs count of 189.4, 76.7, 102.6, and 26.8, respectively. These findings delineate a clear trade-off between accuracy and computational efficiency.
@@
When evaluating overall performance, which includes accuracy, size, and model efficiency, YOLO11m emerges as a consistently top-performing model. It achieves a mAP50-95 of 0.795, an inference time of 2.4ms, a model size of 38.8Mb, and a 67.9 GFLOPs count, as detailed in Figures 11, 13, and Table VI. This is followed by YOLO11l (mAP50-95 of 0.794, inference time of 4.6ms, size of 49Mb, and 86.8 GFLOPs count), and YOLOv10m (mAP50-95 of 0.781, inference time of 2.4ms, size of 32.1Mb, 63.8 GFLOPs count). These results highlight the robustness of these models in detecting traffic signs of various sizes while maintaining short inference times and small model sizes.
On the contrary, YOLOv9e demonstrates poor performance overall, achieving a relatively low mAP50-95 of 0.764, an inference time of 16.1ms, a large model size of 117.5MB, and one of the highest GFLOPs counts at 198.4. Additionally, YOLOv5um, YOLOv8m, and YOLOv8s performed suboptimally, further indicating the inconsistency in performance among earlier model families.
@@
Notably, the YOLO11 and YOLOv12 families significantly outperform other YOLO families in terms of accuracy and computational efficiency. Their models consistently surpass counterparts from the YOLOv3u, YOLOv5u, YOLOv8, YOLOv9, YOLOv10 families, demonstrating their ability to balance precision and speed effectively in the detection of objects with varying sizes.
IV-A2 Africa Wildlife Dataset
TABLE VII: Evaluation results for the Africa Wildlife dataset.
Versions Precision Recall mAP50 mAP50-95
Preprocess
Time
Inference
Time
Postprocess
Time
Total
Time
GFLOPs Size
YOLOv3u 0.75 0.849 0.874 0.781 0.7 8.5 0.4 9.6 207.86 282.4
YOLOV3u tiny 0.845 0.667 0.772 0.682 1.4 0.7 0.3 2.4 24.44 19
YOLOv5un 0.805 0.679 0.749 0.665 0.6 6.6 0.4 7.6 5.65 7.1
YOLOv5us 0.85 0.777 0.827 0.744 0.5 7.8 0.4 8.7 18.58 23.9
YOLOv5um 0.849 0.701 0.83 0.744 1.1 9.5 0.4 11 50.54 64.1
YOLOv5ul 0.831 0.836 0.886 0.799 0.6 9.7 0.4 10.7 106.85 134.9
YOLOv5ux 0.863 0.795 0.867 0.777 1.1 9.8 0.4 11.3 195.2 246.3
YOLOv8n 0.749 0.688 0.777 0.689 0.6 6.8 0.4 7.8 6.55 8.1
YOLOv8s 0.766 0.788 0.806 0.718 0.6 7.8 0.4 8.8 22.59 28.6
YOLOv8m 0.838 0.805 0.845 0.763 1.6 9.1 0.4 11.1 52.12 78.9
YOLOv8l 0.771 0.789 0.853 0.767 0.6 9.2 0.4 10.2 87.77 165
YOLOv8x 0.902 0.744 0.874 0.78 0.6 9.4 0.4 10.4 136.9 257.7
YOLOv9t 0.792 0.748 0.812 0.731 0.5 10 0.4 10.9 4.93 7.7
YOLOv9s 0.763 0.81 0.828 0.75 0.6 11.1 0.4 12.1 15.33 26.8
YOLOv9m 0.864 0.796 0.864 0.784 1 12.1 0.4 13.5 40.98 76.7
YOLOv9c 0.827 0.807 0.852 0.769 1.3 11.6 0.4 13.3 51.8 102.6
YOLOv9e 0.819 0.824 0.854 0.764 0.8 16.1 0.4 17.3 117.5 189.4
YOLOv10n 0.722 0.602 0.722 0.64 1 0.8 0.2 2 5.59 8.3
YOLOv10s 0.823 0.742 0.834 0.744 1.2 1.1 0.2 2.5 15.9 24.7
YOLOv10m 0.834 0.843 0.88 0.781 1.2 2.4 0.2 3.8 32.1 63.8
YOLOv10b 0.836 0.764 0.859 0.765 1 3.1 0.2 4.3 39.7 98.4
YOLOv10l 0.873 0.807 0.866 0.771 1.1 3.8 0.2 5.1 50 126.8
YOLOv10x 0.773 0.854 0.88 0.787 1 6.3 0.2 7.5 61.4 170.4
YOLO11n 0.768 0.695 0.757 0.668 1.2 0.6 0.4 2.2 5.35 6.4
YOLO11s 0.819 0.758 0.838 0.742 1.2 1 0.4 2.6 18.4 21.4
YOLO11m 0.898 0.826 0.893 0.795 1.2 2.4 0.4 4 38.8 67.9
YOLO11l 0.862 0.839 0.889 0.794 1.2 3 0.4 4.6 49 86.8
YOLO11x 0.819 0.816 0.885 0.784 0.9 6.1 0.4 7.4 109 194.8
YOLOv12n 0.883 0.684 0.802 0.708 1.2 4.6 0.4 6.2 5.34 6.4
YOLOv12s 0.761 0.839 0.877 0.78 1.6 4.3 0.5 6.4 18.1 21.4
YOLOv12m 0.868 0.811 0.89 0.79 0.9 4.9 0.4 6.2 39 67.3
YOLOv12l 0.826 0.836 0.877 0.781 0.9 5.5 0.4 6.8 51.2 89.6
YOLOv12x 0.832 0.849 0.889 0.793 1 9.5 0.5 11 114 198.9
@@ IV-A2 Africa Wildlife Dataset
The results in Table VII showcase the performance of the YOLO models on the Africa Wildlife dataset. This dataset contains large object sizes focusing on the ability of YOLO models to predict large objects and their risk of overfitting due to the size of the dataset. The models demonstrate robust accuracy across the board, with the highest-performing models achieving a mAP50-95 ranging from 0.832 to 0.725. This relatively shorter range reflects the effectiveness of the models in detecting and classifying large wildlife objects by maintaining high accuracy.
@@
Figure 14: mAP50 and mAP50-95 YOLO results on Africa wildlife dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score.
@@
Figure 15: Precision vs. Recall based on size results on Africa wildlife dataset, with larger circles indicating larger model sizes.
@@
Figure 16: Total processing time and GFLOPs count results on Africa wildlife dataset.
@@
As illustrated in Figure 14, YOLOv9s demonstrates exceptional performance with a high mAP50-95 of 0.832 and a mAP50 of 0.956, showcasing its robust accuracy across various IoU thresholds. YOLOv9c and YOLOv9t follow closely, with mAP50 scores of 0.96 and 0.948 and mAP50-95 scores of 0.83 and 0.825, respectively. These results highlight the YOLOv9 family’s ability to effectively learn patterns from a small sample of images, making it particularly suited for smaller datasets. In contrast, YOLOv5un, YOLOv10n, and YOLOv3u-tiny show lower mAP50-95 scores of 0.791, 0.786, and 0.725, indicating their limitations in accuracy. The underperformance of larger models like YOLO11x, YOLOv5ux, YOLOv5ul, and YOLOv10l can be attributed to overfitting, especially given the small dataset size.
Precision and Recall:
Figure 15 reveals that YOLO8l and YOLO11l achieve the highest precision and recall, with values of 0.942 and 0.937 for precision, and 0.898 and 0.896 for recall, respectively. Notably, YOLOv8n achieves similar results (0.932 for precision, 0.908 for recall) with a compact size of 6.55Mb, demonstrating its efficiency. In contrast, YOLOv3u and YOLOv5ul exhibit lower precision and recall scores (0.91 and 0.88 for YOLOv3u, 0.916 and 0.881 for YOLOv5ul), despite their larger sizes (204.86Mb for YOLOv3u, 106.85Mb for YOLOv5ul), which may be attributed to overfitting issues.
@@
YOLOv10n, YOLOv8n, and YOLOv3u-tiny are the fastest models, achieving processing times of 2ms and 1.8ms, with GFLOPs counts of 8.2 and 19.1, respectively. The first two models share the same processing speed and GFLOPs count, as showcased in Figure 16. Conversely, YOLOv12x exhibits the slowest processing time at 12.8ms and a GFLOPs count of 198.5, followed by YOLOv9e at 8.5ms and 189.3 GFLOPs count. These results indicate that larger models tend to require more processing time and hardware usage compared to smaller models, emphasizing the trade-off between model size and processing efficiency.
@@
YOLOv9t and YOLOv9s consistently excel across all metrics, delivering high accuracy while maintaining small model sizes, low GFLOPs, and short inference times, as shown in Table VII, and Figures 15 and 16. This demonstrates the robustness of YOLOv9’s smaller models and their effectiveness on small datasets. In contrast, YOLOv12x and YOLO5ux show suboptimal accuracy despite their larger sizes and longer inference times, likely due to overfitting. Most large models underperformed on this dataset, with the exception of YOLOv10x, which benefited from a modern architecture that prevents overfitting.
TABLE VIII: Evaluation results for the ships and vessels dataset.
Versions Precision Recall mAP50 mAP50-95
Preprocess
Time
Inference
Time
Postprocess
Time
Total
Time
Size GFLOPs
YOLOv3u 0.679 0.534 0.519 0.298 0.8 6.2 0.3 7.3 207.86 282.5
YOLOV3u tiny 0.647 0.511 0.489 0.273 1 0.7 0.3 2 24.44 18.9
YOLOv5un 0.635 0.532 0.514 0.298 1.5 0.6 0.3 2.4 5.65 7.2
YOLOv5us 0.653 0.541 0.518 0.299 1.2 0.8 0.3 2.3 18.58 24
YOLOv5um 0.667 0.541 0.526 0.308 0.9 2.1 0.3 3.3 50.54 64
YOLOv5ul 0.654 0.545 0.525 0.305 0.9 3.3 0.3 4.5 106.85 134.8
YOLOv5ux 0.668 0.555 0.531 0.309 0.8 6.7 0.3 7.8 195.2 246.2
YOLOv8n 0.655 0.533 0.515 0.297 1.5 0.5 0.3 2.3 6.55 8
YOLOv8s 0.647 0.545 0.518 0.301 1.1 1 0.3 2.4 22.59 28.5
YOLOv8m 0.669 0.547 0.525 0.302 0.8 2.5 0.3 3.6 52.12 79
YOLOv8l 0.659 0.551 0.526 0.303 0.9 3.9 0.3 5.1 87.77 165
YOLOv8x 0.655 0.55 0.529 0.306 0.8 7.1 0.3 8.2 136.9 257.7
YOLOv9t 0.647 0.516 0.512 0.3 1.4 1.1 0.3 2.8 4.93 7.5
YOLOv9s 0.655 0.552 0.522 0.308 1.4 1.2 0.3 2.9 15.33 26.9
YOLOv9m 0.668 0.551 0.529 0.307 1.1 2.7 0.3 4.1 40.98 76.8
YOLOv9c 0.663 0.547 0.523 0.303 1.2 3.4 0.3 4.9 51.8 102.4
YOLOv9e 0.667 0.537 0.524 0.308 1.1 7.6 0.3 9 117.5 189.5
YOLOv10n 0.584 0.487 0.506 0.31 1.4 0.8 0.2 2.4 5.59 8.2
YOLOv10s 0.586 0.511 0.515 0.319 1.1 1.1 0.2 2.4 15.9 24.4
YOLOv10m 0.588 0.517 0.522 0.322 1 2.4 0.1 3.5 32.1 63.4
YOLOv10b 0.603 0.509 0.523 0.319 1.1 3.2 0.1 4.4 39.7 97.9
YOLOv10l 0.601 0.511 0.522 0.322 1.1 3.8 0.1 5 50 126.3
YOLOv10x 0.6 0.523 0.526 0.321 1 6.3 0.2 7.5 61.4 169.8
YOLO11n 0.574 0.51 0.505 0.311 1.5 0.7 0.3 2.5 5.35 6.3
YOLO11s 0.585 0.535 0.521 0.323 1.3 1 0.3 2.6 18.4 21.3
YOLO11m 0.588 0.541 0.53 0.325 1 2.4 0.3 3.7 38.8 67.6
YOLO11l 0.596 0.531 0.528 0.325 1.1 3 0.4 4.5 49 86.6
YOLO11x 0.596 0.538 0.529 0.327 0.8 6.1 0.3 7.2 109 194.4
YOLOv12n 0.571 0.505 0.505 0.307 1.1 4.7 0.4 6.2 5.34 6.3
YOLOv12s 0.58 0.529 0.516 0.319 1.6 5 0.4 7 18.1 21.2
YOLOv12m 0.584 0.543 0.525 0.325 1.1 4.6 0.4 6.1 39 67.1
YOLOv12l 0.588 0.532 0.522 0.323 2.6 5.2 0.4 8.2 51.2 88.5
YOLOv12x 0.586 0.538 0.52 0.321 1.2 9.3 0.4 10.9 114 198.5
Figure 17: mAP50 and mAP50-95 YOLO results on ships and vessel dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score.
Figure 18: Precision vs. Recall based on size results on ships and vessels dataset, with larger circles indicating larger model sizes.
Figure 19: Total processing time and GFLOPs count results on ships and vessels dataset.
IV-A3 Ships and Vessels Dataset
Table VIII presents the performance of YOLO models on the Ships and Vessels dataset, a large dataset featuring tiny objects with varying orientations. Overall, the models demonstrated moderate effectiveness in detecting ships and vessels, with mAP50-95 ranging from 0.273 to 0.327. This performance suggests that YOLO algorithms may face challenges in accurately detecting smaller objects with different orientations.
@@ IV-A3 Ships and Vessels Dataset
The disparity between mAP50-95 and mAP50, illustrated in Figure 17, underscores the challenges YOLO models face with higher IoU thresholds when detecting small objects. Additionally, YOLO models struggle with detecting objects of varying rotations. Among the models, YOLO11x achieved the highest accuracy, with a mAP50 of 0.529 and a mAP50-95 of 0.327, closely followed by YOLO11l, YOLO11m, and YOLO11s, which recorded mAP50 values of 0.529, 0.528, and 0.53, and mAP50-95 values of 0.327, 0.325, and 0.325, respectively. These results highlight the robustness of the YOLO11 family in detecting small and tiny objects. In contrast, YOLOv3u-tiny, YOLOv8n, YOLOv3u, and YOLOv5n exhibited the lowest accuracy, with mAP50 scores of 0.489, 0.515, 0.519, and 0.514, and mAP50-95 scores of 0.273, 0.297, 0.298, and 0.298, respectively. This suggests the outdated YOLOv3u architecture and the potential underfitting of smaller models with large size of datasets.
Precision and Recall:
Figure 18 indicates that YOLOv5ux outperformed other models, achieving a precision of 0.668 and a recall of 0.555. It was closely followed by YOLOv9m (precision of 0.668, recall of 0.551) and YOLOv8m (precision of 0.669, recall of 0.525), both of which are significantly smaller in size (40.98 Mb for YOLOv9m and 52.12 Mb for YOLOv8m). In contrast, YOLOv12n and YOLO11n exhibited lower performance, with precisions of 0.571 and 0.574 and recalls of 0.505 and 0.51, respectively, likely due to underfitting issues. Generally, YOLOv12 and YOLO11 models tended to produce false positives, reflected in their low precision and high recall. Meanwhile, YOLOv10 underperformed in both precision and recall, despite being one of the newest models in the YOLO family.
@@
As illustrated in Figure 19, YOLOv3u-tiny achieved the fastest processing time at 2 ms, closely followed by YOLOv8n and YOLOv5un, both recording 2.3 ms. YOLOv10 and YOLO11 models also excelled in speed, with YOLOv10n and YOLO11n achieving rapid inference times of 2.4 ms and 2.5 ms, along with GFLOPs counts of 8.2 and 6.3, respectively. In contrast, YOLOv12x exhibited the lowest speed, with 10.9 ms inference time and 198.5 GFLOPs count, highlighting the shortcomings of the newest YOLO family.
@@
The results in Table VIII and Figures 17, 18, and 19 demonstrate that YOLO11s and YOLOv10s excelled in accuracy while maintaining compact sizes, low GFLOPs, and quick processing times. In contrast, YOLOv3u, YOLOv8x, and YOLOv8l fell short of expectations despite their larger sizes and longer processing times. These findings highlight the robustness and reliability of the YOLO11 family, particularly in improving the YOLO family’s performance in detecting small and tiny objects while ensuring efficient processing. Additionally, the results reveal the underperformance of YOLOv9 models when faced with large datasets and small objects, despite their modern architecture.
IV-B Discussion
Based on the performance of the models across the three datasets, we ranked them by accuracy, speed, GFLOps count, and size, as shown in Table IX to facilitate a comprehensive evaluation. For accuracy, the mAP50-95 metric was employed due to its capacity to assess models across a range of IoU thresholds, thus providing a detailed insight into each model’s performance. For speed, models were sorted based on the total processing time, which encompasses preprocessing, inference, and postprocessing durations. The rankings range from Rank 1, indicating the highest performance, to Rank 28, denoting the lowest, with the respective rankings highlighted in bold within the table.
TABLE IX: Overall ranking of YOLO algorithms

### v4

@@ 4.1.1 Traffic Signs Dataset
Figure 12 presents a comparative analysis of the YOLO algorithms’ performance on the Traffic Signs dataset, evaluated based on accuracy, computational efficiency, and model size. The Traffic Signs dataset is a medium-sized dataset with varied object sizes, making it favorable for benchmarking.
@@
As illustrated in Figure 13, YOLOv5ul demonstrates the highest accuracy, achieving a mAP50 of 0.866 and a mAP50-95 of 0.799. This is followed by YOLO11m with a mAP50-95 of 0.795 and YOLO11l with a mAP50-95 of 0.794. In contrast, YOLOv10n exhibits the lowest precision, with a mAP50 of 0.722 and a mAP50-95 of 0.64, closely followed by YOLOv5un with a mAP50-95 of 0.665.
Figure 13: mAP50 and mAP50-95 YOLO results on traffic signs dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score
Figure 14: Precision vs. Recall based on size on traffic signs dataset, with larger circles indicating larger model sizes
Figure 15: Total processing time and GFLOPs count results on traffic signs dataset
@@
Regarding computational efficiency, YOLOv10n is the most efficient, with a processing time of 2ms per image and a GFLOPs count of 8.3, as shown in Figure 15. YOLO11n closely trails this at 2.2ms with a 6.4 GFLOPs count, and YOLOv3u tiny with a processing time of 2.4ms and a GFLOPs count of 19, making it relatively computationally inefficient compared to the other fast models. However, the data indicates that YOLOv9e, YOLOv9m, YOLOv9c, and YOLOv9s are the least efficient, with inference times of 16.1ms, 12.1ms, 11.6ms, and 11.1ms, and GFLOPs count of 189.4, 76.7, 102.6, and 26.8, respectively. These findings delineate a clear trade-off between accuracy and computational efficiency.
@@
When evaluating overall performance, which includes accuracy, size, and model efficiency, YOLO11m emerges as a consistently top-performing model as detailed in Figures 12, 13, 14, and 15. This is followed by YOLO11l and YOLOv10m showcasing the robustness of these models. On the contrary, YOLOv9e demonstrates poor performance overall in terms of accuracy, speed, GFLOPs, and model size. Additionally, YOLOv5um, YOLOv8m, and YOLOv8s performed suboptimally, further indicating the inconsistency in performance among earlier model families. Notably, the YOLO11 and YOLOv12 families significantly outperform other YOLO families regarding accuracy and computational efficiency. Their models consistently surpass counterparts from the YOLOv3u, YOLOv5u, YOLOv8, YOLOv9, and YOLOv10 families, demonstrating their ability to balance precision and speed effectively in the detection of objects with varying sizes.
4.1.2 Africa Wildlife Dataset
Figure 16: Evaluation results for the Africa wildlife dataset.
@@ 4.1.2 Africa Wildlife Dataset
The results in Figure 16 showcase the performance of the YOLO models on the Africa Wildlife dataset. This dataset contains large object sizes, focusing on the ability of YOLO models to predict large objects and their risk of overfitting due to the size of the dataset.
@@
Figure 17: mAP50 and mAP50-95 YOLO results on Africa wildlife dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score
@@
Figure 18: Precision vs. Recall based on size results on Africa wildlife dataset, with larger circles indicating larger model sizes
@@
Figure 19: Total processing time and GFLOPs count results on Africa wildlife dataset
@@
As illustrated in Figure 17, YOLOv9s demonstrates exceptional performance with a high mAP50-95 of 0.832 and a mAP50 of 0.956, showcasing its robust accuracy across various IoU thresholds. YOLOv9c and YOLOv9t follow closely, with mAP50 scores of 0.96 and 0.948 and mAP50-95 scores of 0.83 and 0.825, respectively. These results highlight the YOLOv9 family’s ability to effectively learn patterns from a small sample of images, making it particularly suited for smaller datasets. In contrast, YOLOv5un, YOLOv10n, and YOLOv3u tiny show lower mAP50-95 scores of 0.791, 0.786, and 0.725, indicating their limitations in accuracy. The underperformance of larger models like YOLO11x, YOLOv5ux, YOLOv5ul, and YOLOv10l can be attributed to overfitting, especially given the small dataset size.
@@
YOLOv10n, YOLOv8n, and YOLOv3u tiny are the fastest models, achieving processing times of 2ms and 1.8ms, with GFLOPs counts of 8.2 and 19.1, respectively. The first two models share the same processing speed and GFLOPs count, as showcased in Figure 19. Conversely, YOLOv12x exhibits the slowest processing time at 12.8 ms and a GFLOPs count of 198.5, followed by YOLOv9e at 8.5 ms and 189.3 GFLOPs count. These results indicate that larger models tend to require more processing time and hardware usage compared to smaller models, emphasizing the trade-off between model size and processing efficiency.
@@
YOLOv9t and YOLOv9s consistently excel across all metrics, delivering high accuracy while maintaining small model sizes, low GFLOPs, and short inference times, as shown in Figures 16, 18, 18, and 19. This demonstrates the robustness of YOLOv9’s smaller models and their effectiveness on small datasets. In contrast, YOLOv12x and YOLO5ux show suboptimal accuracy despite their larger sizes and longer inference times, likely due to overfitting. Most large models underperformed on this dataset, with the exception of YOLOv10x, which benefited from a modern architecture that prevents overfitting.
Figure 20: Evaluation results for the ships and vessels dataset
Figure 21: mAP50 and mAP50-95 YOLO results on ships and vessel dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score
Figure 22: Precision vs. Recall based on size results on ships and vessels dataset, with larger circles indicating larger model sizes
Figure 23: Total processing time and GFLOPs count results on ships and vessels dataset
4.1.3 Ships and Vessels Dataset
Figure 20 presents the performance of YOLO models on the Ships and Vessels dataset, a large dataset featuring tiny objects with varying orientations.
@@ 4.1.3 Ships and Vessels Dataset
The disparity between mAP50-95 and mAP50, illustrated in Figure 21, underscores the challenges YOLO models face with higher IoU thresholds when detecting small objects. Additionally, YOLO models struggle with detecting objects of varying rotations. Among the models, YOLO11x achieved the highest accuracy, with a mAP50 of 0.529 and a mAP50-95 of 0.327, closely followed by YOLO11l, YOLO11m, and YOLO11s, which recorded mAP50 values of 0.529, 0.528, and 0.53, and mAP50-95 values of 0.327, 0.325, and 0.325, respectively. In contrast, YOLOv3u-tiny, YOLOv8n, YOLOv3u, and YOLOv5n exhibited the lowest accuracy, with mAP50 scores of 0.489, 0.515, 0.519, and 0.514, and mAP50-95 scores of 0.273, 0.297, 0.298, and 0.298, respectively. This suggests the outdated YOLOv3u architecture and the potential underfitting of smaller models with large datasets.
@@
As illustrated in Figure 23, YOLOv3u tiny achieved the fastest processing time at 2 ms, closely followed by YOLOv8n and YOLOv5un, recording 2.3 ms. YOLOv10 and YOLO11 models also excelled in speed, with YOLOv10n and YOLO11n achieving rapid inference times of 2.4 ms and 2.5 ms, along with GFLOPs counts of 8.2 and 6.3, respectively. In contrast, YOLOv12x exhibited the lowest speed, with 10.9 ms inference time and 198.5 GFLOPs count, highlighting the shortcomings of the newest YOLO family.
@@
The results in Figures 20, 21, 22, and 23 demonstrate that YOLO11s and YOLOv10s excelled in accuracy while maintaining compact sizes, low GFLOPs, and quick processing times. In contrast, YOLOv3u, YOLOv8x, and YOLOv8l fell short of expectations despite their larger sizes and longer processing times. These findings highlight the robustness and reliability of the YOLO11 family, particularly in improving the YOLO family’s performance in detecting small and tiny objects while ensuring efficient processing. Additionally, the results reveal the underperformance of YOLOv9 models when faced with large datasets and small objects despite their modern architecture.
5 Discussion
Models were ranked by accuracy, speed, GFLOPs count, and size across three datasets, as shown in Figure 24. To ensure fair comparisons, models were categorized by scale. YOLOv3u tiny was classified as small, YOLOv9t as nano, YOLOv9c as medium, YOLOv9c and YOLOv10b as large, and YOLOv9e as extra-large. Additionally, family-wide performance was analyzed by averaging rankings across all scales to compare overall performance as seen in Table 4.
Accuracy was measured using mAP50-95 for a comprehensive performance assessment, while speed rankings were based on total processing time . Rankings range from Rank 1 (best) to the lowest performer, with results highlighted in bold. This approach ensures a fair evaluation of architectural improvements and computational trade-offs across model sizes and families.
Figure 24: Overall ranking of YOLO algorithms based on size
Table 4: Overall ranking of YOLO algorithms based on family